OpenStudy (anonymous):

An ice cube has mass 9.0g and is added to a cup of coffee. The coffee's intial temp. is 90.0 deg. Celsius and the cup contains 120.0 g of liquid. Assume the specific heat capacity of coffee is the same as that of water. The heat of fusion of ice (associated with ice melting) is 6.0kJ/mol. Find the temp. of coffee after ice melts. How do we get rid of the /mol under the kJ? Also, may I get a hint please?

OpenStudy (anonymous):

are we using deg C or deg F? but i'll assume its deg C to get rid of the mol, just find the molecular mass of ice which is H2O, 18g/mol so therefore, specific latent heat of fusion of ice = 6.0kJ/18g = 333.3J/g latent heat of fusion of ice = (6.0kJ/18)*9 = 3.0kJ ok i am not too sure what value you are using for specific heat capacity of water, but lets just use 4.20J/g therefore 3.0kJ = 120*c*(change in temperature) = 504* change in temperature change in temperature = 3000/504 = 5.952 K Hence, final temperature = 90.0 - 5.952 = 84.0 deg C